OpenStudy (anonymous):

5/4 in simplest form

OpenStudy (frizz):

Well that's simple, take five divided by four.

OpenStudy (anonymous):

thanks lol..

OpenStudy (frizz):

No problem!

OpenStudy (frozen15):

5/4 in simplest form 5/4 5/4 = 1.25 as a decimal form 5/4 = 1.25 in 2 decimal places 5/4 = 1.3 to the nearest tenth 5/4 = 1.25 to the nearest hundredth 5/4 = 1.25 to the nearest thousandth heres more examples incase

OpenStudy (texaschic101):

it is in simplest form...it cannot be reduced

OpenStudy (frozen15):

yup that's why when I wrote this , it said simplest form and under it , It had shown 5/4 still its whole

OpenStudy (frozen15):

does anyone get a medal?